6x^2 + x - 2 = 0
I am just learning this...Help!

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

6x^2 + x - 2 = 0

Factoras follows:
(3x+2)(2x-1)=0
Since the product is zerot one
of these factors must be zero.
3x+2 =0 or 2x-1=0
If 3x+2=0
Then 3x=-2
x=-2/3
--------
If 2x-1=0
Then 2x=1
x=1/2
So, x=-2/3 or x=1/2
Cheers,
Stan H.

Answer by checkley71(8403)

6X^2+X-2=0 YOU NEED TO FIND TWO FACTORS OF 6 (1,6 OR 2.3) THAT WHEN MULTIPLIED BY THE FACTORS OF -2 (1,-2 OR -1,2) EQUALS 1.
THUS
(3X+2)(2X-1)=0
NOW SOLVE BOTH OF THESE FACTORS BY SETTING EACH EQUAL T0 0
3X+2=0
3X=-2
X=-2/3
AND
2X-1=0
2X=1
X=1/2
PROOF IS REPLACING X BY -2/3 & THEN BY 1/2